# OpenVIS: Open-vocabulary Video Instance Segmentation

### v2

@@ Abstract
Open-vocabulary Video Instance Segmentation (OpenVIS) can simultaneously detect, segment, and track arbitrary object categories in a video, without being constrained to categories seen during training. In this work, we propose an OpenVIS framework called InstFormer that achieves powerful open vocabulary capability through lightweight fine-tuning on a limited-category labeled dataset. Specifically, InstFormer comes in three steps a) Open-world Mask Proposal: we utilize a query-based transformer, which is encouraged to propose all potential object instances, to obtain class-agnostic instance masks; b) Open-vocabulary Instance Representation and Classification: we propose InstCLIP, adapted from pre-trained CLIP with Instance Guidance Attention. InstCLIP generates the instance token capable of representing each open-vocabulary instance. These instance tokens not only enable open-vocabulary classification for multiple instances with a single CLIP forward pass but have also been proven effective for subsequent open-vocabulary instance tracking. c) Rollout Association: we introduce a class-agnostic rollout tracker to predict rollout tokens from the tracking tokens of previous frames to enable open-vocabulary instance association across frames in the video. The experimental results demonstrate the proposed InstFormer achieve state-of-the-art capabilities on a comprehensive OpenVIS evaluation benchmark, while also achieves competitive performance in fully supervised VIS task.
Keywords:
Open-vocabulary Video Instance Segmentation Instance Guidance CLIP
1 Introduction
@@
Video understanding [3, 19, 13, 34, 46, 25] is a challenging yet significant computer vision task that requires specialized algorithms and techniques, surpassing the difficulty of image understanding. It has a wide range of applications spanning various real-world scenarios, such as surveillance, robotics, and autonomous driving. To achieve a more thorough understanding to target video, Video Instance Segmentation (VIS) [41] has been proposed, which can simultaneously detect, segment, and track instances in a given video, becoming a new research hotspot. Recently, the introduction of query-based detectors from image perception domain has contributed to further advancements in VIS. Despite the remarkable advancements, the current VIS models possess an inherent limitation. They can only segment objects within the boundaries of their training data, meaning they are unable to identify objects beyond the categories present in the training set. As a result, their understanding of target videos remains constrained. Furthermore, every time the well-trained model is required to identify novel categories, it necessitates additional training using data that includes objects from new categories. This process demands substantial human annotations, resulting in significant investments of time and resources.
@@
Figure 1: Visualization of Open-vocabulary Video Instance Segmentation. OpenVIS simultaneously segments, detects, and tracks arbitrary objects in a video according to their corresponding text description. As illustrated above, the proposed InstFormer can accurately identify various objects based on their respective category names in a video, irrespective of whether the category is included in the training set.
To address the limitation, we investigate a novel computer vision task called Open-vocabulary Video Instance Segmentation (OpenVIS). This task focuses on detecting, segmenting, and tracking instances in videos based on the category names of target objects, regardless of whether those categories have been seen during the training stage. Although recent pre-trained Vision-Language Models (VLMs) [30, 43] have shown promising results in zero-shot classification, there are still several challenges that need to be addressed in order to achieve open-vocabulary video instance segmentation. Firstly, OpenVIS requires a mask proposal network that can segment as many instances as possible in a video, since every instance has the potential to be identified in this task. Secondly, directly inputting the masked proposals to pre-trained VLMs, such as CLIP [30], is not an optimal choice for zero-shot instence classification, since 1) the distribution gap between training images (i.e., natural images) and test images (i.e., masked images) and 2) inefficient inference process that it necessitates performing CLIP vision-encoder computations $N$ times for $N$ instance proposals per frame (as shown in Fig. 2 (a)). Thirdly, the training process of current instance trackers with fixed-category datasets in Video Instance Segmentation presents a significant challenge when it comes to tracking open-vocabulary instances.
To solve the above issues, we propose a novel open-vocabulary video instance segmentation framework, InstFormer. Specifically, we first employ a query-based mask proposal network to generate class-agnostic instance masks, and introduce a margin instance contrastive loss to the network, with the goal of promoting the proposal of a greater number of distinct instances within a given video. Then, we propose the InstCLIP, a variant of CLIP enhanced with the proposed Instance Guidance Attention, which directs instance tokens to attend different instance regions by the generated guidance attention mask, enabling efficient open-vocabulary classification for multiple mask proposals. By leveraging Instance Guidance Attention, InstCLIP enables a single-forward open-vocabulary classification for multiple proposals within a single image, eliminating the need to repeatedly input the masked image for each instance. Moreover, InstCLIP has a natural byproduct, i.e., the instance tokens, which naturally aligns with instance queries from the mask proposal network, further enhancing the association of open-vocabulary instances across frames. Specifically, instance tokens can be concatenated with instance proposal queries to form the tracking tokens for open-vocabulary instance association. Finally, we introduce a rollout tracker in our proposed framework. This tracker utilizes a simple yet history-aware RNN layer to predict rollout tokens based on previous tracking tokens. By employing class-independent temporal contrastive learning, the rollout tracker enables open-vocabulary instance association.
@@
To facilitate research on this novel task, we propose an evaluation benchmark that utilizes readily available datasets to thoroughly assess the performance. In our benchmark, the OpenVIS model will be trained with a limited number of categories, and subsequently tested on a large number of categories. Specifically, we evaluate the proposed model on YouTube-VIS, BURST, LVVIS, and UVO datasets, encompassing a large number of novel categories, to comprehensively assess its diverse capacities. However, the training process only see the data of YouTube-VIS, which comprises only 40 categories.
Figure 2: Open-vocabulary instance classification paradigm comparison in OpenVIS. Our InstCLIP is not only more efficient compared to inputting masked images $N$ times for $N$ instances, but the resulting instance token can also be utilized for tracking.
The experimental results demonstrate the proposed InstFormer achieves state-of-the-art capabilities in OpenVIS and competitive performance in fully supervised VIS. Our contributions can be summarized as follows:
  • We propose InstFormer to explore the novel computer vision task, i.e., Open-vocabulary Video Instance Segmentation (OpenVIS), and sugeest a comprehensive evaluation benchmark.
@@
  • We present InstCLIP, an open-vocabulary instance representation and classification network designed to embed each open-vocabulary instance with an instance token. The resulting instance tokens not only enable efficient open-vocabulary classification for multiple instances but also prove effective in subsequent open-vocabulary instance tracking.
  • We propose a class-independent rollout tracker that utilizes temporal contrastive learning to associate instances of any category across consecutive frames in a video.
2 Related Work
2.1 Video Instance Segmentation
@@
Video instance segmentation was initially introduced by [41]. There are two main prevailing paradigms for VIS: offline approaches and online approaches. Offline approaches such as VisTR [35], Mask2Former-VIS [7], and SeqFormer [36], take the whole video as input and utilize instance queries to predict the instance sequence of the entire video or clip in a single step. Despite the high performance on popular datasets, the requirement of the whole video limits the application of offline methods, especially for long video and on-going video scenarios. Online approaches like MaskTrack R-CNN [41], MaskProp [2], MinVIS [22], IDOL [37], GenVIS [18], and DVIS [44] independently process each frame to obtain all the mask proposals and their corresponding categories and then track these instance proposals by post-processing methods. Our approach also builds on this way which provides us with more flexibility and convenience in predicting potential mask proposals and using InstCLIP for instance classification.
2.2 Vision-Language Models
@@
Vision-language models aim to bridge the gap between visual and textual modalities, gaining significant attention due to their impressive performance in visual representation learning. Recently, benefiting from the large-scale training data, pre-trained VLMs like CLIP [30] and FLIP [43] have further shown strong zero-shot object recognition capability. For example, after training on 400 million image-text paired data, CLIP (VIT-L as visual encoder) can achieve 76.2% zero-shot accuracy on ImageNet, encouraging a series of downstream vision tasks to involving these pre-trained models, such as classification [30, 23], captioning [21], retrieval [29], and segmentation [39]. Despite VLMs enabling powerful open-vocabulary capability to downstream tasks, directly involving VLMs in some tasks is non-trivial, e.g., the original VLMs are trained on non-masked images, and as a consequence, their performance naturally declines with the masked input images in video instance segmentation [26]. Additionally, the computational cost of running $N$ times VLMs’ vision encoder for $N$ instance within a frame is impractical.
2.3 Open-Vocabulary Segmentation
@@
Open vocabulary segmentation is proposed by ZS3Net [5], which segments objects by their corresponding text description, which may contain categories not have been seen during the training stage. The mainstream works, such as ZSSeg [39], ZegFormer [11], and OVSeg [26], utilize a two-stage framework to achieve open vocabulary segmentation, i.e., first extracting all potential proposals without classes from an image, then calculating the similarity between the visual features of proposals and text features of potential text descriptions to identify the categories. Recent concurrent work SAN [38] and DeOP [15] begin addressing the impractical computation cost of multiple passes of VLMs vision encoder in above methods, similar to the InstCLIP in our framework. However, compared to the above works, InstCLIP only needs a lightweight finetuning without changing VLM pre-trained weights and provides instance token for follow-by video-level instance association.
3 Setting
3.1 Problem Formulation
@@ 3.2 Evaluation Benchmark
  • Open-world Property: We leverage the exhaustively annotated UVO dataset [33] to evaluate the open-world mask proposal ability. The UVO dataset provides an average of 13.52 instances annotated per video. Compared to the only 1.68 objects in YouTube-VIS, UVO is naturally a suitable dataset for measuring open-world property.
  • Zero-shot Property: We utilize the category-rich BURST dataset [1] to evaluate the zero-shot instance classification property. The objects in BURST involve 482 categories, with 78 common categories from COCO [27] and 404 uncommon categories, which can be regard as unseen categories. The uncommon-404 categories is an ideal dataset to measure the zero-shot property. Additionally, a latest dataset LVVIS [32] novel set contains 555 unseen categories for our setting, so we also evaluate it.
@@
Following [41], we utilize the Average Precision (i.e., AP) and Average Recall (i.e., AR) at the video level as the main metrics. Additionally, our OpenVIS model is only trained on YouTube-VIS (a widely-used VIS dataset comprising 40 categories). This ensures that the categories present in the training data are small-scale subsets of those found in the test data. More discussion and analysis of the evaluation benchmark can be found in Supplementary.
4 Method
In this section, we introduce the proposed InstFormer in details (illustrated in Fig. 3). Specifically, we first introduce how we fine-tuned an open-world mask proposal network with the objective of predicting class-agnostic mask proposals of all potential objects in Sec. 4.1. Then, we propose InstCLIP for efficient open-vocabulary instance representation and classification in Sec. 4.2. Within the InstCLIP, instances are embedded with instance tokens, which not only facilitate open-vocabulary classification for multiple instances with a single CLIP forward pass but also significantly benefit the follow-up open-vocabulary tracking as well. Finally, we propose a class-independent rollout tracker in Sec. 4.3 that leverages temporal contrastive learning to perform online open-vocabulary instance association across frames with a simple yet history-aware RNN layer.
Figure 3: Overview of the proposed InstFormer framework for open-vocabulary video instance segmentation. (a) Open-world Mask Proposal (Sec. 4.1): Generate class-agnostic instance masks with a query-based transformer, which is encouraged to propose all potential object instances. (b) Open-vocabulary Instance Representation and Classification (Sec. 4.2): Perform open-vocabulary classification for the multiple instances with the proposed InstCLIP. Meanwhile, the instance tokens, which can facilitate follow-up instance tracking, are also acquired in this process. (c) Rollout Association (Sec. 4.3): Associate instances of any categories across frames with the proposed class-independent rollout tracker driven by the temporal contrastive learning, which consists of a simple yet history-aware RNN layer.
4.1 Open-world Mask Proposal
An open-world proposed mask proposal network needs to propose as many distinct instances as possible to meet the flexible needs of open-world perception, as each instance has the possibility of being selected to identify by end user. To achieve this goal, we first adopted a universal image segmentation model Mask2Former [8] as the mask proposal network, predicting $N$ class-agnostic masks $M_{t}=\{m_{t}^{i}\}_{i=1}^{N}\in[0,1]^{N\times H\times W}$ and their corresponding instance queries $Q_{t}=\{q_{t}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ for each frame $F_{t}\in\mathbb{R}^{3\times H\times W}$ of a video:
@@ 4.1 Open-world Mask Proposal
4.2 Open-vocabulary Instance Representation and Classification
[…]
$X_{t}^{l}=\mathrm{InstAttn}(X_{t}^{l-1},\mathcal{M})=\mathrm{softmax}(W^{q}X_{t}^{l-1}\cdot W^{k}X_{t}^{l-1}+\mathcal{M})\cdot W^{v}X_{t}^{l-1},$ (3)
[…]
where $W^{q},W^{k},W^{v}$ are weights of query, key, and value projection layer, respectively. $X_{t}^{l-1}$ consists of vision tokens $V_{t}\in\mathbb{R}^{\frac{H}{32}\times\frac{W}{32}\times C}$ from image patch embedding, $N$ initial instance tokens $I^{l}\in\mathbb{R}^{N\times C}$, and a register token $R^{l}\in\mathbb{R}^{1\times C}$. $P$ is the number of vision tokens. The register token, inspired by [10], is a token permitted to attend to all vision tokens. It plays the role of collecting low-informative feature, which helps obtain cleaner attention maps from instance tokens to vision tokens.
@@ 4.2 Open-vocabulary Instance Representation and Classification
where $E\in\mathbb{R}^{K\times C}$ is the vocabulary embeddings of $K$ categories. InstCLIP is designed with the principle of minimizing modifications from CLIP, to fully unleash the zero-shot capability of the pre-trained CLIP. Only the linear projections for the query and value of the attention layer are adjusted using the parameter-efficient fine-tuning approach LoRA [20] during training, while almost parameters of CLIP remain frozen.
4.3 Rollout Association
Association with Instance Tracking Tokens.
Within the proposed InstFormer framework, we have instance queries $Q_{t}$ from the query-based mask proposal network and instance tokens $I_{t}$ naturally aligned with instance queries from the instance-motivated InstCLIP. The former, capable of generating class-agnostic mask proposals for open-world instances, has been demonstrated in MinVIS [22] can associate instances between frames with simple similarity matching. The latter, representing instances with CLIP features having the capability of open vocabulary, also proves to be particularly helpful for tracking open-vocabulary instances in Sec. 5.3. Hence, we obtain the instance tracking token by combining them: $T^{Tr}_{t}=\mathrm{Cat}(Q_{t},I_{t})\in\mathbb{R}^{N\times 2C}$. These open-vocabulary instances can be associated across frames based on the tracking tokens through the assignment of applying the Hungarian matching on the similarity score $S_{ij}=\mathrm{cos}(T^{Tr}_{t},T^{Tr}_{t-1})$, without any additional parameters being trained in this tracking process.
Rollout Tracker.
The above association mechanism makes an assumption not entirely reasonable: the tracking tokens for objects between different frames should be very similar. Clearly, instances in the video undergo varying degrees of appearance changes, so expecting their tracking tokens to remain consistent is not practical. Also, relying solely on the previous frame for tracking is non-robust, as an error in one frame can result in lasting errors in subsequent frames.
Here we propose a concise yet effective rollout tracker that predicts rollout association tokens $T^{A}_{t}\in\mathbb{R}^{N\times C}$ of the current frame from tracking tokens of previous frames, to replace the tracking token of the previous frame in association process:
@@ 4.3 Rollout Association
where $\mathcal{R}$ denotes the rollout tracker implemented by a single RNN layer and $h_{t-1}$ is the hidden state of RNN remaining instance temporal information. The rollout association token can be considered as a prediction of the current frame’s tracking token derived from past frames, which is more suitable for matching with the actual tracking tokens of the current frame. The rollout tracker is trained with a temporal contrastive loss, which pulls tokens of the same instance closer and pushes tokens of different instances farther away:
$\mathcal{L}_{TC}=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathrm{CE}(\mathrm{cos}(T^{A}_{t}(i),T^{Tr}_{t-1}(j)),\mathbbm{1}_{[i=j]}),$ (6)
where $\mathbbm{1}_{[i=j]}\in\{0,1\}$ is an indicator function evaluating to $1$ if $i=j$. This temporal contrastive loss is independent of categories, allowing the rollout tracker to handle open-vocabulary instances. Meanwhile, the introduced historical information in this history-aware tracker also addresses the limitation of robustness associated with tracking solely based on the previous frame.
5 Experiment
5.1 Implementation Details
Model Architecture.
We regard a COCO [27]-pretrained Mask2Former [8] as our mask proposal network. By default, the transformer decoder has 100 queries, with a dimension of 256 for the query embedding and transformer decoder. For InstCLIP, we select a ViT-B/32 of CLIP [30] as foundation vision transformer. The number of instance tokens of InstCLIP is also set to 100, aligning with the 100 instance queries of the mask proposal network. We initialize the instance tokens and register token using CLIP’s learned class tokens. The text encoder is a 12-layer transformer, the same as that in CLIP. For input prompts, we ensemble 14 prompts (e.g., “a photo of a {category name}”) from [26] to boost zero-shot classification ability.
Training.
InstFormer is trained using a two-stage approach and CLIP weights are frozen during the entire training. In first stage, the open-world mask proposal network and InstCLIP (LoRA adapter) are trained for 6k iterations with $\mathcal{L}_{SC}$ and instance segmentation loss. Subsequently, we train the rollout tracker in second stage, with all other weights frozen, using $\mathcal{L}_{TC}$ for an additional 600 iterations. The whole training is done on 8 V100 GPUs for 3 hours.
Baselines.
To better assess the performance of the proposed InstFormer framework, we introduce several baselines for comparison, as shown in Tab. 1. For fully-supervised methods, we provide STCN Tracker [1], Box Tracker [1] and MinVIS [22]. Both the first two methods ultilize the Mask-RCNN [17] but with different tracking strategies as in [1]. MinVIS [22] is an advanced VIS model in fully-supervised VIS task. All of them are trained on full BURST dataset with all 482 categories. For open-vocabulary methods, we employ three approaches: Detic-SORT, Detic-OWTB, and OV2Seg [32]. The first two methods utilize the open-vocabulary detector Detic [45], paired with the classical multi-object tracker SORT [4] and the state-of-the-art open-world tracker OWTB [28], respectively. OV2Seg introduces the CLIP text encoder and a momentum-updated query for tracking, also achieving open-vocabulary video instance segmentation. These baselines may be trained with different datasets. We provide their corresponding training datasets and category numbers in Tab. 1. L-1203 represents the entire LVIS [14] dataset with all 1203 categories. L-866 indicates the LVIS subset with 866 frequent categories. B and Y denote BURST [1] with 482 categories and YouTube-VIS [41] with 40 categories, respectively. C is the COCO [27] dataset with 80 categories, which may be used to pretrain some modules. And for clarity, we provide the proportion of novel categories during inference for each approach and setting in the table in gray(higher means more challenging).
5.2 Main Results
Overall Performance.
[…]

### v3

@@ Abstract
Open-vocabulary Video Instance Segmentation (OpenVIS) can simultaneously detect, segment, and track arbitrary object categories in a video, without being constrained to categories seen during training. In this work, we propose InstFormer, a carefully designed framework for the OpenVIS task that achieves powerful open-vocabulary capabilities through lightweight fine-tuning with limited-category data. InstFormer begins with the open-world mask proposal network, encouraged to propose all potential instance class-agnostic masks by the contrastive instance margin loss. Next, we introduce InstCLIP, adapted from pre-trained CLIP with Instance Guidance Attention, which encodes open-vocabulary instance tokens efficiently. These instance tokens not only enable open-vocabulary classification but also offer strong universal tracking capabilities. Furthermore, to prevent the tracking module from being constrained by the training data with limited categories, we propose the universal rollout association, which transforms the tracking problem into predicting the next frame’s instance tracking token. The experimental results demonstrate the proposed InstFormer achieve state-of-the-art capabilities on a comprehensive OpenVIS evaluation benchmark, while also achieves competitive performance in fully supervised VIS task.
Introduction
@@
Video understanding (Bertasius, Wang, and Torresani 2021; Hong et al. 2022; Guo et al. 2022; Wang et al. 2022; Zhou et al. 2023; Li et al. 2023) is a challenging yet significant computer vision task that requires specialized algorithms and techniques, surpassing the difficulty of image understanding. To achieve a more thorough understanding, Video Instance Segmentation (VIS) (Yang, Fan, and Xu 2019) has been proposed, which can simultaneously detect, segment, and track instances in a given video, becoming a new research hotspot. Despite significant progress, current VIS models possess an inherent limitation. They can only segment objects within the boundaries of their training data, meaning they are unable to identify objects beyond the categories present in the training set. Consequently, their video understanding remains restricted. Moreover, identifying new categories requires retraining with additional annotated data, leading to substantial time and resource investment. To address the limitation, we investigate a novel computer vision task called Open-vocabulary Video Instance Segmentation (OpenVIS). This task focuses on detecting, segmenting, and tracking instances in videos based on the category names of target objects, regardless of whether those categories have been seen during the training stage.
@@
Figure 1: Visualization of Open-vocabulary Video Instance Segmentation. OpenVIS simultaneously segments, detects, and tracks arbitrary objects in a video according to their corresponding text description. The proposed InstFormer can accurately identify various objects based on their respective category names in a video, irrespective of whether the category is included in the training set.
Although recent pre-trained Vision-Language Models (VLMs) (Radford et al. 2021; Yao et al. 2021) have shown promising results in zero-shot classification and provide good foundation for open-vocabulary video instance segmentation, significant challenges still remain in leveraging these static, image-level VLMs for this video, instance-level task. To eliminate this gap, we propose InstFormer, a carefully designed framework tailored for the OpenVIS task that achieves robust open-vocabulary capabilities through lightweight fine-tuning on a limited-category labeled dataset. Firstly, since any object, rather than fixed categories, might be selected for identification by the end user, InstFormer first performs the open-world mask proposal by incorporating a margin instance contrastive loss into a query-based mask proposal network to generates class-agnostic instance masks, with the goal of proposing as many distinct instances within a given video as possible to meet the flexible needs of open-world perception. Secondly, obtaining open-vocabulary representations of instances to enable classification and tracking across frames is non-trivial. Leveraging the zero-shot capabilities of pre-trained VLMs like CLIP (Radford et al. 2021) by directly inputting masked instance images is suboptimal and inefficient for video tasks due to the domain gap between pre-training on natural images and testing on masked images, as well as the need to run the VLM’s vision encoder multiple times per frame. So we propose the InstCLIP, a variant of CLIP adapted with the proposed Instance Guidance Attention, which directs instance tokens to attend different instance regions simultaneously by the generated guidance according to multiple mask proposals. These instance tokens not only enable open-vocabulary classification but also offer strong universal tracking capabilities. Thirdly, the training process of current instance trackers with fixed-category datasets in video instance segmentation presents a significant challenge when it comes to tracking open-vocabulary instances. To address this issue, we propose Universal Rollout Association, where the rollout tracker is trained to predict instance tokens of the next frame to achieve tracking. The rollout tracker is implemented with a simple yet history-aware RNN layer, predicting instance tokens for the next frame based on previous tracking tokens. This prediction training is independent of categories, enabling the rollout tracker to handle open-vocabulary task.
@@
To facilitate research on this novel task, we propose an evaluation benchmark that utilizes readily available datasets to thoroughly assess the performance. In our benchmark, the OpenVIS model will be trained with a limited number of categories, and subsequently tested on a large number of categories. Specifically, we evaluate the proposed model on YouTube-VIS, BURST, LVVIS, and UVO datasets, encompassing a large number of novel categories, to comprehensively assess its diverse capacities. However, the training process only see the data of YouTube-VIS, which comprises only 40 categories. The experimental results demonstrate the proposed InstFormer achieves state-of-the-art capabilities in OpenVIS and competitive performance in fully supervised VIS. This indicates that InstFormer retains most of VLM’s zero-shot capabilities while optimizing for specific domains, providing a sound solution for scenarios needing both extreme domain performance and generalization. Our contributions can be summarized as follows:
  • We propose the InstFormer framework, which achieves open-vocabulary capabilities through lightweight fine-tuning on limited-category data, to explore the novel OpenVIS task and introduce a comprehensive evaluation benchmark.
@@
  • We present InstCLIP, designed to embed each open-vocabulary instance with an instance token. The resulting instance tokens not only enable efficient open-vocabulary classification for multiple instances but also prove effective in subsequent open-vocabulary instance tracking.
  • We propose universal rollout association, which achieves tracking by training the tracker to predict instance tokens of the next frame, to overcome the limitations of trackers trained on fixed-category data that struggle to generalize to open-vocabulary instances.
Related Work
Video Instance Segmentation
@@
There are two main prevailing paradigms for Video Instance Segmentation (Yang, Fan, and Xu 2019): offline approaches and online approaches. Offline approaches such as VisTR (Wang et al. 2021b), Mask2Former-VIS (Cheng et al. 2021), and SeqFormer (Wu et al. 2021), take the whole video as input and utilize instance queries to predict the instance sequence of the entire video or clip in a single step. Despite the high performance on popular datasets, the requirement of the whole video limits the application of offline methods, especially for long video and on-going video scenarios. Online approaches like MaskTrack RCNN (Yang, Fan, and Xu 2019), MaskProp (Bertasius and Torresani 2020), MinVIS (Huang et al. 2022), IDOL (Wu et al. 2022a), and DVIS (Zhang et al. 2023) independently process each frame to obtain all the mask proposals and their corresponding categories and then track these instance proposals by post-processing methods. Our approach also builds on this way which provides us with more flexibility and convenience in predicting potential mask proposals and using InstCLIP for instance classification.
Vision-Language Models
@@
Vision-language models aim to bridge the gap between visual and textual modalities, gaining significant attention due to their impressive performance in visual representation learning. Recently, benefiting from the large-scale training data, pre-trained VLMs like CLIP (Radford et al. 2021) and FLIP (Yao et al. 2021) have further shown strong zero-shot object recognition capability. For example, after training on 400 million image-text paired data, CLIP (VIT-L as visual encoder) can achieve 76.2% zero-shot accuracy on ImageNet, encouraging a series of downstream vision tasks to involving these pre-trained models, such as classification (Radford et al. 2021; Huang, Chu, and Wei 2022), captioning (Hu et al. 2022), retrieval (Liu et al. 2021), and segmentation (Xu et al. 2021). Despite VLMs enabling powerful open-vocabulary capability to downstream tasks, directly involving VLMs in some tasks is non-trivial, e.g., the original VLMs are trained on non-masked images, and as a consequence, their performance naturally declines with the masked input images in video instance segmentation (Liang et al. 2022). Additionally, the computational cost of running $N$ times VLMs’ vision encoder for $N$ instance within a frame is impractical.
Open-Vocabulary Segmentation
@@
Open vocabulary segmentation is proposed by ZS3Net (Bucher et al. 2019), which segments objects by their corresponding text description, which may contain categories not have been seen during the training stage. The mainstream works, such as ZSSeg (Xu et al. 2021), ZegFormer (Ding et al. 2022), and OVSeg (Liang et al. 2022), utilize a two-stage framework to achieve open vocabulary segmentation, i.e., first extracting all potential proposals without classes from an image, then calculating the similarity between the visual features of proposals and text features of potential text descriptions to identify the categories. Recent concurrent work SAN (Xu et al. 2023) and DeOP (Han et al. 2023) begin addressing the impractical computation cost of multiple passes of VLMs vision encoder in above methods, similar to the InstCLIP in our framework. However, compared to the above works, InstCLIP only needs a lightweight finetuning without changing VLM pre-trained weights and provides instance token for follow-by video-level instance association.
Setting
Problem Formulation
@@
  • Open-world Property: We leverage the exhaustively annotated UVO dataset (Wang et al. 2021a) to evaluate the open-world mask proposal ability. The UVO dataset provides an average of 13.52 instances annotated per video. Compared to the only 1.68 objects in YouTube-VIS, UVO is naturally a suitable dataset for measuring open-world property.
  • Zero-shot Property: We utilize the category-rich BURST dataset (Athar et al. 2023) to evaluate the zero-shot instance classification property. The objects in BURST involve 482 categories, with 78 common categories from COCO (Lin et al. 2014) and 404 uncommon categories, which can be regard as unseen categories. The uncommon-404 categories is an ideal dataset to measure the zero-shot property. Additionally, a latest dataset LVVIS (Wang et al. 2023) novel set contains 555 unseen categories for our setting, so we also evaluate it.
@@
Following (Yang, Fan, and Xu 2019), we utilize the Average Precision (i.e., AP) and Average Recall (i.e., AR) at the video level as the main metrics. Additionally, our OpenVIS model is only trained on YouTube-VIS (a widely-used VIS dataset comprising 40 categories). This ensures that the categories present in the training data are small-scale subsets of those found in the test data. More discussion and analysis of the evaluation benchmark can be found in Supplementary.
Figure 2: Overview of the proposed InstFormer framework for OpenVIS. (a) Open-world Mask Proposal: Generate class-agnostic instance masks with a query-based transformer, which is encouraged to propose all potential object instances. (b) Open-vocabulary Instance Representation: InstCLIP embeds open-vocabulary instance tokens using Instance Guidance Attention efficiently. These tokens enable open-vocabulary instance classification and provide robust open-vocabulary tracking capabilities. (c) Universal Rollout Association: Associate instances of any category across frames with the proposed universal rollout tracker, which is trained to predict the instance tracking tokens of the next frame, termed the rollout token.
Method
In this section, we detail how we bridge the gap between static, image-level VLMs and the video, instance-level demands of the OpenVIS task, leading to the InstFormer, a carefully designed framework tailored for OpenVIS that achieves open-vocabulary capabilities through lightweight fine-tuning on limited-category labeled data (as illustrated in Fig. 2).
Open-world Mask Proposal
An open-world proposed mask proposal network needs to propose as many distinct instances as possible to meet the flexible needs of open-world perception, as each instance has the possibility of being selected to identify by end user. To achieve this goal, we first adopted a query-based image segmentation model Mask2Former (Cheng et al. 2022) as the mask proposal network, predicting $N$ class-agnostic masks $M_{t}=\{m_{t}^{i}\}_{i=1}^{N}\in[0,1]^{N\times H\times W}$ and their corresponding instance queries $Q_{t}=\{q_{t}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ for each frame $F_{t}\in\mathbb{R}^{3\times H\times W}$ of a video:
@@
Open-vocabulary Instance Representation
Leveraging pre-trained VLMs like CLIP (Radford et al. 2021) for zero-shot capabilities by directly inputting masked instance images is suboptimal and inefficient for real-time video tasks due to the domain gap between pre-training on natural images and testing on masked images, as well as the need to run CLIP vision encoder $N$ times per frame. The proposed InstCLIP efficiently represents each instance with an instance token, enabling open-vocabulary instance classification and providing strong open-vocabulary tracking capabilities. Specifically, InstCLIP is a Vision Transformer (Dosovitskiy et al. 2020) adapted from the pre-trained CLIP vision encoder, consisting of $L$ Instance Guidance Attention layers. We generate attention masks from mask proposals for Instance Guidance Attention to guide $N$ instance tokens to embed $N$ instances in a single forward pass through the encoder. Instance Guidance Attention takes as input the concatenated tokens $X_{t}^{l-1}\in\mathbb{R}^{1+N+P}$ from the previous attention layer and the guidance attention mask $\mathcal{M}\in\mathbb{R}^{(1+N+P)\times(1+N+P)}$:
$X_{t}^{l}$ $=\mathrm{InstAttn}(X_{t}^{l-1},\mathcal{M})$ (3)
$\quad=\mathrm{softmax}(W^{q}X_{t}^{l-1}\cdot W^{k}X_{t}^{l-1}+\mathcal{M})\cdot W^{v}X_{t}^{l-1},$
where $W^{q},W^{k},W^{v}$ are weights of query, key, and value projection layer, respectively. $X_{t}^{l-1}$ consists of vision tokens $V_{t}\in\mathbb{R}^{\frac{H}{32}\times\frac{W}{32}\times C}$ from image patch embedding, $N$ initial instance tokens $I^{l}\in\mathbb{R}^{N\times C}$, and a register token $R^{l}\in\mathbb{R}^{1\times C}$. $P$ is the number of vision tokens. The register token, inspired by (Darcet et al. 2023), is a token permitted to attend to all vision tokens. It plays the role of collecting low-informative feature, which helps obtain cleaner attention maps from instance tokens to vision tokens.
@@
where $E\in\mathbb{R}^{K\times C}$ is the vocabulary embeddings of $K$ categories. InstCLIP is designed with the principle of minimizing modifications from CLIP, to fully unleash the zero-shot capability of the pre-trained CLIP. Only the linear projections for the query and value of the attention layer are adjusted using the parameter-efficient fine-tuning approach LoRA (Hu et al. 2021) during training, while almost parameters of CLIP remain frozen.
Universal Rollout Association
To prevent trackers optimized on closed-set data from failing to generalize to open-vocabulary instance tracking, the proposed Universal Rollout Association fully leverages the open-vocabulary characteristics of instance tokens and transforms the tracking problem into predicting the next frame’s instance tracking token for training.
Instance Tracking Tokens.
We form universal instance tracking tokens $T^{Tr}_{t}$ by combining the instance tokens $I_{t}$ from InstCLIP with the instance queries $Q_{t}$ from the proposal network to handle open-vocabulary tracking: $T^{Tr}_{t}=\mathrm{Concat}(I_{t},Q_{t})\in\mathbb{R}^{N\times 2C}$. These tokens and queries are naturally aligned in a token/query-based architecture. The former, which leverages CLIP features with zero-shot capabilities, proves particularly effective for tracking open-vocabulary instances in subsequent experiments. The latter, characterized by its ability to generate class-agnostic mask proposals, has also been demonstrated in MinVIS (Huang et al. 2022) to distinguish instances between frames.
Rollout Association.
And to prevent the tracker from being constrained by fixed-category object data, we reframe the tracking problem by training the tracker to predict the instance tracking token for the next frame. When associating instances in frame-t with those from previous frames, the rollout tracker predicts the instance tracking token for frame-t based on the tokens from history frames. This predicted instance tracking token, referred to as the rollout association token, is denoted as $T^{A}_{t}\in\mathbb{R}^{N\times C}$:
@@
where $\mathcal{R}$ denotes the concise yet effective rollout tracker implemented by a single RNN layer and $h_{t-1}$ is the hidden state of RNN remaining instance temporal information. Finally, by comparing the rollout association token for frame t with the actual instance tracking token using Hungarian matching on the similarity score $S_{ij}=\mathrm{cos}(T^{A}_{t},T^{Tr}_{t})$, the instance association for frame-t can be completed. The rollout tracker is trained with the loss:
$\mathcal{L}_{T}=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathrm{CE}(\mathrm{cos}(T^{A}_{t}(i),T^{Tr}_{t}(j)),\mathbbm{1}_{[i=j]}),$ (6)
where $\mathbbm{1}_{[i=j]}\in\{0,1\}$ is an indicator function evaluating to $1$ if $i=j$. This process is independent of categories, allowing the rollout tracker to handle open-vocabulary instances. Meanwhile, the incorporation of historical information in this history-aware tracker enhances robustness compared to tracking based solely on the previous frame.
Experiment
Method OV Training BURST LVVISnovel